OpenStudy (anonymous):

\[x^2-y^2=a,2xy=b,\text{find }x^2+y^2\]

OpenStudy (anonymous):

in terms of a and b

OpenStudy (asnaseer):

subtract the second expression from the first one and see if you can spot something

OpenStudy (anonymous):

\[x^2-(\frac{b}{2x})^2=x^2-\frac{b^2}{4x^2}=a\] \[\text{let }x^2=k\] \[4k^2-4ak-b^2=0\] \[k=\frac{4a\pm \sqrt{16a^2+16b}}{8}=\frac{a\pm \sqrt{a^2+b}}{2}\] so \[y=\frac{b^2}{4\frac{a\pm \sqrt{a^2+b}}{2}}=\frac{b^2}{2 a\pm \sqrt{a^2+b}}\] \[x^2+y^2=\frac{a\pm \sqrt{a^2+b}}{2}+\frac{b^2}{2 a\pm \sqrt{a^2+b}}\]

OpenStudy (anonymous):

\[x^2-y^2-2xy=(x-y)^2\]

OpenStudy (asnaseer):

yup - then factorise the first expression

OpenStudy (anonymous):

wait somethings wrong

OpenStudy (anonymous):

\[x^2+y^2+2xy=(x+y)^2=a+b\]

OpenStudy (asnaseer):

where did you get that from?

OpenStudy (anonymous):

oh its when i open this \[(x+y)^2=x^2+2xy+y^2\]

OpenStudy (asnaseer):

I mean where did you get the right-hand-side from (a+b)?

OpenStudy (anonymous):

ohh mistake

OpenStudy (asnaseer):

:)

OpenStudy (anonymous):

this question is actually a follow up of this http://openstudy.com/study#/updates/51fd18d1e4b05ca7841f512c

OpenStudy (asnaseer):

ok - but if you follow my suggestion then you should be able to solve this

OpenStudy (anonymous):

\[x^2-y^2-2xy=?\]

OpenStudy (anonymous):

a-b

OpenStudy (asnaseer):

i.e. you know:\[x^2-y^2-2xy=a-b\]therefore:\[(x-y)^2=a-b\]therefore:\[x-y=\sqrt{a-b}\tag{1}\]and from the first expression you get:\[x^2-y^2=a\]therefore:\[(x+y)(x-y)=a\]then use expression (1) to find x+y

OpenStudy (asnaseer):

then use:\[(x+y)^2=x^2+y^2+2xy\]

OpenStudy (anonymous):

\[ (x+y)^2=\frac{a^2}{(\sqrt{a-b})^2}\] so \[x^2+y^2=\frac{a^2}{a-b}-b\]

OpenStudy (anonymous):

\[\frac{a^2+b^2-ab}{a-b}\]

OpenStudy (asnaseer):

looks right

OpenStudy (anonymous):

if we go back to our original problem \[(x^2+y^2)^2=\frac{(a^2+b^2-ab)^2}{(a-b)^2}\] is required

OpenStudy (asnaseer):

when you say "original problem" do you mean the one referred to by the link you posted above?

OpenStudy (anonymous):

yes the one on the link,for case \[a=\frac{ac+bd}{c^2+d^2},b=\frac{bc+ad}{c^2+d^2}\]

OpenStudy (asnaseer):

let me take a look at the original to see if there is an easier way to get to the solution...

OpenStudy (anonymous):

yes the one on the link,for case \[a=\frac{ac+bd}{c^2+d^2},b=\frac{bc+ad}{c^2+d^2}\]

OpenStudy (anonymous):

\[b=\frac{bc-ad}{c^2+d^2}\] we need to show that \[(x^2+y^2)^2=\frac{a^2+b^2}{c^2+d^2}\]

OpenStudy (shubhamsrg):

let x = sqrta cost and y = i sqrta sint so 2ia sint cost = b ia sin2t = b you have to find value of a(cos^2 t - sin^2 t) or acos2t = asqrt(1- sin^2 (2t)) just substitute value of sin2t

OpenStudy (anonymous):

\[\text{let x} = \sqrt a \cos t\] and \[y = i \sqrt a \sin t\] so \[2ia \sin t \cos t = b\] \[ia \sin2t = b\] you have to find value of \[a(\cos^2 t -\sin^2 t) ,OR a\cos2t\] \[= a\sqrt{(1- \sin^2 (2t)) }\] just substitute value of sin2t i juss wanted to see wats happening clearly,thanks @shelby1985

OpenStudy (anonymous):

@shubhamsrg

OpenStudy (shubhamsrg):

:)

OpenStudy (anonymous):

\[a\sqrt{(1-\frac{b^2}{-a^2}})=a\sqrt{\frac{a^2+b^2}{a^2}}=\sqrt{a^2+b^2}\]

OpenStudy (anonymous):

is this final?

OpenStudy (shubhamsrg):

according to me,yes

OpenStudy (anonymous):

thanks @asnaseer @shubhamsrg @DLS

OpenStudy (asnaseer):

yw :)

OpenStudy (anonymous):

on this answer wat happened to c^2+d^2 denoinator

OpenStudy (asnaseer):

In your original problem you got to:\[\begin{align} x^2-y^2&=\frac{ac+bd}{c^2+d^2}\tag{1}\\ 2xy&=\frac{bc-ad}{c^2+d^2}\tag{2}\\ &\text{using the substitutions suggested by shubhamsrg, we get:}\\ x&=\sqrt{\frac{ac+bd}{c^2+d^2}}\cos(t)\tag{3}\\ y&=i\sqrt{\frac{ac+bd}{c^2+d^2}}\sin(t)\tag{4}\\ &\text{this leads to:}\\ \sin(2t)&=\frac{ad-bc}{ac+bd}i\tag{5}\\ \therefore x^2+y^2&=\frac{ac+bd}{c^2+d^2}\sqrt{1-\sin^2(2t))}\\ &=\frac{ac+bd}{c^2+d^2}\sqrt{1+\frac{(ad-bc)^2}{(ac+bd)^2}}\\ &=\frac{ac+bd}{c^2+d^2}\sqrt{\frac{(ac+bd)^2+(ad-bc)^2}{(ac+bd)^2}}\\ &=\frac{\sqrt{(ac+bd)^2+(ad-bc)^2}}{c^2+d^2}\\ &=\frac{\sqrt{a^2c^2+2abcd+b^2d^2+a^2d^2-2abcd+b^2c^2}}{c^2+d^2}\\ &=\frac{\sqrt{a^2c^2+b^2d^2+a^2d^2+b^2c^2}}{c^2+d^2}\\ &=\frac{\sqrt{a^2(c^2+d^2)+b^2(c^2+d^2)}}{c^2+d^2}\\ &=\frac{\sqrt{(a^2+b^2)(c^2+d^2)}}{c^2+d^2}\\ &=\sqrt{\frac{a^2+b^2}{c^2+d^2}}\\ \therefore (x^2+y^2)^2&=\frac{a^2+b^2}{c^2+d^2}\\ \end{align}\]

OpenStudy (anonymous):

can someone give asneer a medal pls

OpenStudy (anonymous):

thank you again

OpenStudy (asnaseer):

no need for medal? I am just here to help and learn :)

OpenStudy (anonymous):

thank you soo much ,i was wondering wat is keeping you typing for so loooooooong

OpenStudy (asnaseer):

:) - and thx Joe for the medal :)